Justify your steps using the commutative, associative, and distributive properties when necessary
y+4+2x=3y
I tried to help my child but have no idea please help

Answer by rfer(16322)

y+4+2x=3y
subtract y from both sides
y-y+4+2x=3y-y
4+2x=2y
subtract 4 from both sides
4-4+2x=2y-4
2x=2y-4
divide both sides by 2
2x/2=(2y-4)/2
x=y-2